OpenStudy (openstudy9):

Need some help with this coordinate geometry problem! Will medal+fan! Circle K lies In the x-y coordinate plane, has its center at (-2,5), and passes through the point (1,1). Which of the following points does circle K NOT pass through? A. (1,9) B.(-5,9) C.(-5,1) D.(-2,0) E. (0,5)

Nnesha (nnesha):

general form of circle \[\large\rm (x-h)^2+(y-k)^2=r^2\] where (h,k) is the center and (x,y) the point and r is the radius so first substitute (x,y) for given point and (h,k) for (-2,5) then solve for r

OpenStudy (openstudy9):

(1+2)^2+(1-5)^2

OpenStudy (openstudy9):

radius is 5, right?

Nnesha (nnesha):

correct now substitute (h,k) with center and r with 5 then test each option substitute (x,y) for each option and see which one give you the true statement

Nnesha (nnesha):

actually u have to find which one NOT true

OpenStudy (openstudy9):

The answer is E

OpenStudy (openstudy9):

Right?

Nnesha (nnesha):

why do you think its E ? post the work :=))

OpenStudy (openstudy9):

Becuase: (0+2)^2+(5-5)^2=4, BUT THE RADIUS SQUARED NEEDS TO BE 25

OpenStudy (openstudy9):

right?

Nnesha (nnesha):

YES lel

Nnesha (nnesha):

good work!